# Distilling the Essence: Efficient Reasoning Distillation via Sequence Truncation

### v1

@@ 3.3 Methods
Section-wise Supervision (RQ1). The supervision comprises three sections: P, CoT, and A and we vary which sections are involved in backpropagation. The six conditions are A, P+A, CoT, CoT+A, P+CoT, and P+CoT+A (full-length with all sections included). Following the training setup of previous work (Mukherjee et al., 2023; Mitra et al., 2023; Jiang et al., 2023; Gu et al., 2024; Shi et al., 2024; Shen et al., 2025; Li et al., 2025a), we apply a loss mask to all tokens outside the supervised region before computing the objective, while still feeding the entire P+CoT+A sequence through the model in the forward pass. If a section is not included, tokens within that section are conditioned but not involved in loss computation.
@@
Sequence-Length Scaling Behavior. First, we explore how performance scales with the amount of tokens trained. Unlike loss masking, which still incurs forward-pass costs, we examine the performance when tokens are removed entirely. We introduce a dynamic truncation protocol, Lead-Span Proportion (LSP). For each example with tokenized inputs $\mathbf{x}$ and labels $\mathbf{y}$, LSP retains the first $p\in\{0.1,0.2,\dots,1.0\}$ fraction of tokens from both $\mathbf{x}$ and $\mathbf{y}$, and removes the rightmost (1p)(1-p) fraction before training. The truncated tokens, therefore, never enter the model: they are neither seen in the forward pass nor included in the loss. This protocol preserves example-specific variability in kept length, avoids systematic exclusion of answer tokens that can arise under a global max-length cutoff, and maintains a diversity of sequence lengths.
Budget Location Ablation. Second, to determine where the budget should be spent, we fix the training budget at $50\%$ of the tokens (a choice guided by performance saturation, compute efficiency, and experimental symmetry, as detailed in Section 4.2). We compare two conditions: Left 50%, which keeps the first half of tokens and drops the rest, and Right 50%, which keeps the final half and drops the beginning. This ablation keeps the compute budget fixed while shifting the kept segment of tokens, enabling a clean test of whether early tokens carry richer learning signal than later ones. Better performance in the Left 50% setting would be consistent with the hypothesis that useful signal concentrates early in the trajectory.
@@ 4.1 Section-wise Supervision (RQ1)
As can be seen in Figure 1, a consistent pattern emerges across benchmarks and teacher-student pairs: settings that supervise with CoT consistently outperform those that do not.
On both benchmarks, answer-only supervision (A) yields the lowest accuracy, and extending the loss to the prompt (P+A) brings only modest gains. Once CoT tokens receive supervision, whether alone (CoT), paired with the answer (CoT+A), paired with the prompt (P+CoT), or combined with both (P+CoT+A), performance improves substantially, with only minor differences among these CoT-inclusive variants.
@@
Overall, the results indicate that where supervision is applied along the trajectory matters more than how many tokens receive loss. Supervising the rationale tokens is crucial, while spreading loss to the prompt and answer offers only modest, sometimes negligible, additional benefits. Thus, full-token supervision does not reliably outperform more selective objectives that focus on the reasoning process itself.
To understand the limited gains from supervising P and A, we manually examined the data and observed that the CoT often encapsulates information from both the prompt and the answer as illustrated in Figure 2. Typically, the CoT begins by paraphrasing the question to contextualize the problem, then proceeds with reflection-based reasoning steps before deriving the final answer.
@@
To verify whether P and A are commonly covered by the CoT in our training datasets, we conduct a linguistic analysis on 100 randomly sampled training examples using GPT-5.1 as a judge. We prompt the model to verify that: (1) all substantive information in P and A is already present in the CoT; and (2) P and A do not introduce any new substantive statements absent from the CoT. This ensures that, semantically, the CoT entails both P and A. We also checked whether the final answer in A is equivalent to the final answer derived in the CoT.
@@
Figure 5: First half of tokens tends to contain denser useful information than the second half, in line with stronger performance when training on the left $50\%$ of tokens. Two complete examples are illustrated in Figure 7 and Figure 8 in Appendix A.1.
@@
Figure 3 shows that across all settings, LSP exhibits similar scaling behavior. They show nearly monotonic accuracy improvement as LSP increases and then saturates. The flattening near the right edge suggests those trimmed tokens are not required for effective distillation, implying substantial compute savings with minimal performance drop.
Across all three datasets, the 32B-8B setting is consistently the top performer, and the ranking is largely determined by student size: 8B > 4B > 1.7B > 0.6B. For a fixed student size, curves from different teachers track closely. For the 4B student model, it performs comparably and sometimes better when distilling from 8B rather than 32B model, which echoes the findings of Li et al. (2025b) that smaller models do not necessarily learn better from much larger teachers.
For small students (0.6B and 1.7B), training with LSP does not reliably beat the baselines. Their curves hew close to the dotted baselines across p, suggesting limited capacity to absorb additional supervision. The benefits of LSP mainly materialize once the student has enough scale (4B and 8B). Therefore, in later experiments, these two small students will not be involved.
@@
Deciding the most cost-effective budget. To anchor the budget allocation, we select $p=0.5$ as our target. This decision is driven by three factors: performance saturation, computational efficiency, and experimental rigor. First, we identified the most cost-effective point (“knee”) in the accuracy-vs-LSP curves with  Satopaa et al. (2011) [^12] https://github.com/arvkevi/kneed, as shown in Table 3. We observe that across all settings, the knee consistently occurs at $p\leq 0.6$, with the majority at or below $p=0.5$. Second, as detailed in Section 4.3, $p=0.5$ serves as a critical boundary before computational costs begin to rise drastically. Third, selecting $p=0.5$ enables a clean, symmetrical ablation between the Left 50% and Right 50%. A budget of $p>0.5$ would necessitate overlapping tokens.
As seen in Table 4, across 3 teacher-student pairs, 2 datasets, and 2 benchmarks, training with the first half of tokens retains at least $\approx 86\%$ of the accuracy when compared to training with all tokens, with retention being as high as $\approx 99\%$ for AIME25 with 8B-4B. We further validate these findings on the SkyT1 dataset. The results on SkyT1 (detailed in Appendix A.4) align with those on OpenThoughts and Bespoke, confirming that the performance retention is a generalizable phenomenon across training datasets produced by different models (DeepSeek-R1 and QwQ-32B-Preview).
Openthoughts Bespoke
Pairs AIME24 AIME25 AIME24 AIME25
32B-4B 0.962 0.974 0.980 0.933
8B-4B 0.981 0.989 0.858 0.876
32B-8B 0.960 0.976 0.924 0.872
@@
Table 4: The $\%$ of performance retained by training with just the first half of the tokens of each example in Openthoughts and Bespoke dataset, compared to training with full-length examples.
@@
Figure 6: GPU memory usage, time spent for each training step, and FLOPs as functions of Lead-Span Proportion (LSP). Due to the linear flops behavior, only selected LSPs ($0.5,1.0)$ for some teacher-student pairs are profiled.
@@
Why Early Tokens Suffice? The fact that the first $50\%$ of tokens yields similar performance compared to full-length raises a mechanistic question: does this segment contain a complete reasoning process? To investigate, we analyze where the first derivation of the correct answer appears inside the CoT, regarding it as the effective endpoint of the core reasoning process. For each of $100$ randomly sampled training examples, we prompt GPT-5.1 to identify the substring corresponding to the earliest derivation of the correct answer. We then map this substring to its token index to compute its relative position within the CoT. For Openthoughts and Bespoke, the average relative positions of the first derivation of correct answer are at $50.5\%$ and $47.5\%$ of the CoT, respectively. In terms of the full sequence, these are at approximately $43.3\%$ and $42.0\%$ of the total length. Notably, all these values are less than $50\%$, meaning that the first derivation of the correct answer is, on average, fully included within the left-half training budget. We illustrate this phenomenon in Figure 5.
@@
Qualitative analysis of the content preceding the first derivation shows that this segment contains question understanding, initial ideation, several rounds of self-reflections, and the final answer. This concentration of critical information explains why the first $50\%$ is sufficient to transfer core reasoning capabilities. The remaining tokens are then largely dominated by verification, minor algebraic clean-up, and re-statements of an already-correct answer, which appear less informative for distillation than the earlier reasoning steps. It explains observed little performance gap between half and full-length training. An annotated example is included as Figure 8 in Appendix A.1.
@@
We operationalize self-reflection by counting occurrences of lexical cues such as “wait” and “rethink” (along with related variants; the full keyword list is provided in Appendix A.2.1) that appear before the index of the first correct derivation. For Openthoughts and Bespoke, there are on average $2.2$ and $2.4$ times of self-reflection before first derivation of correct answer, respectively.
@@
Ablation Studies Having established that the left $50\%$ often encapsulates the core reasoning process, we use the Left vs. Right ablation to verify that the efficiency gains are specific to this self-contained initial segment, rather than a general property of reducing the token count. We compare training with the Left $50\%$ against the Right $50\%$. As shown in Figure 4, across all datasets and teacher–student pairs, training with the left $50\%$ of the sequence consistently and significantly outperforms training with the right $50\%$. This is further supported by our loss analysis (Appendix A.3), which reveals that the loss on the left 50% is comparable to the full sequence, whereas the loss on the right 50% is significantly higher, indicating that the initial segment captures the bulk of the learnable signal.
@@ 4.3 Compute Efficiency
Computationally, our $50\%$ training protocol yields substantial resource savings, as illustrated in Figure 6. We observe a critical inflection point at LSP =0.5=0.5 for both memory usage and training time, beyond which costs escalate rapidly due to the quadratic complexity of attention mechanisms.
@@
Training Speed. Training step time follows a similar trajectory. While larger students generally require more time, we also observe that teacher size imposes a distinct overhead; for the same 4B student, distilling from a 32B teacher is consistently slower than from an 8B teacher. By capping the sequence at $50\%$, we achieve a roughly $2\times$ reduction in FLOPs, confirming that $p=0.5$ is a practical sweet spot for resource-constrained distillation.
@@ 5 Conclusion
In this work, we demonstrate that the reasoning trace itself is the dominant carrier of supervisory signal in knowledge distillation. By systematically varying the supervised sections of training examples, we found that models trained on the reasoning trace alone achieve performance comparable to full-sequence supervision. Moreover, our investigation into token budget allocation reveals that critical reasoning behaviors are concentrated in early tokens; truncating examples to just their first $50\%$ of tokens retains $\approx 94\%$ of downstream accuracy on math tasks while reducing training time and memory usage by $\approx 50\%$. This establishes sequence truncation as a powerful efficiency lever, positioning partial-sequence training as a cost-effective strategy for scaling reasoning distillation.
@@ A.2.2 Results
We present the results of our linguistic analysis on the three datasets in Table 6.
@@
P $\subseteq$ CoT A $\subseteq$ CoT Ans Match Pos. Self-Refl.
OpenThoughts 99 89 100 50.5% 2.23
Bespoke 97 93 99 47.5% 2.42
SkyT1 90 73 91 66.8% 1.91
Table 6: Linguistic analysis results on 100 sampled examples per dataset. Columns indicate the percentage of examples where Prompt (P) and Answer (A) are fully covered by the CoT, whether the final answer matches, the average relative token position of the first correct answer derivation, and the average count of self-reflection keywords preceding that derivation.
OpenThoughts and Bespoke.
Semantic Redundancy. The results confirm that for Openthoughts and Bespoke, the CoT almost always contains all substantive information from the Prompt ($97\%$) and the Answer ($89\%$). Furthermore, the final answer derived in the CoT is practically always equivalent to the ground-truth label ($99\%$). This explains why supervising P and A provides little additional signal: the student model can learn the same information purely from the CoT.
Reasoning Concentration. The “Pos.” column shows the average relative position where the correct answer is first derived within the CoT: $50.5\%$ for OpenThoughts, $47.5\%$ for Bespoke. Using the dataset-specific token distribution statistics from Table 2, we calculate the position of the first derivation relative to the full sequence ($P$) as:
$P$ $=Share_{\text{prompt}}+P_{\text{deriv}}\times Share_{\text{cot}}$
$P_{\text{openthought}}$ $=4.7\%+50.5\%\times 76.6\%=43.3\%$
$P_{\text{bespoke}}$ $=5.6\%+47.5\%\times 76.6\%=42.0\%$
These values fall well within the first $50\%$ of the sequence, supporting our finding that the first half of the training sequence contains the core reasoning process for these datasets. The consistent presence of self-reflection keywords before this derivation point ($2.23$ and $2.42$ per example for OpenThoughts and Bespoke, respectively) suggests that this early segment captures critical thinking and error correction, making it effective for distillation even when later tokens are truncated.
@@
Semantic Redundancy. The results for SkyT1 show lower redundancy compared to the other two datasets: Prompt coverage is 90% and Answer coverage is 73% (compared to $97\%$ and $89\%$ for the others). This suggests that for SkyT1, new information or reasoning is relatively more likely to be introduced outside the CoT.
Reasoning Concentration. For SkyT1, the average relative position of the first derivation of the correct answer within CoT is 66.8% which translates to position relative to full sequence as:
$P_{\text{skyt1}}$ $=6.1\%+66.8\%\times 71.7\%=54.0\%$
Analysis of Dataset Characteristics. As seen in Table  2, the distribution of each section of DeepSeek-R1 synthesized datasets (Openthoughts and Bespoke) and QwQ synthesized dataset (SkyT1) are significantly different: the relative length of CoT is shorter in SkyT1 $($$71.7\%-76.6\%)/76.6\%\approx-6.5\%$) while the answer section is significantly longer ($(22.1\%-18.8\%)/18.8\%+(22.1\%-17.8\%)/17.8\%]/2\approx 20.9\%$).
We observe an inconsistent formatting phenomenon in SkyT1, specifically regarding the placement of <begin_of_thought> and <end_of_thought> tokens (which are replaced by <think> and </think>). In contrast, OpenThoughts and Bespoke maintain consistent placement. Out of 8000 training examples in SkyT1, 149 ($\approx 1.86\%$) of the examples do not include exactly one pair of <think> and </think>. On the other hand, all training examples in Openthoughts and Bespoke include exactly one pair of them. For example, one of the SkyT1 examples with misplaced </think> and one additional <think> exhibits the following structure:
[…]
Additionally, SkyT1 sometimes deviates from the pattern where the answer section serves merely as a summary of the CoT as observed in Openthoughts and Bespoke. Instead, SkyT1 is more likely to introduce new reasoning steps or derive new ideas within the answer section (i.e., later tokens). For instance, the reasoning process is not always fully contained between the start and end thinking tokens. This could be the main reason why the value of A $\subseteq$ CoT of SkyT1 (73%) is much lower than that of Openthoughts (89%) and Bespoke (93%). We observe an example where the CoT derives a final answer (e.g. 47) at the end of CoT section, but the answer section continues deriving and reaches a different final answer (e.g. 54) as can be seen as follows:

### v2

@@ 3.3 Methods
Section-wise Supervision (RQ1). The supervision comprises three sections: P, CoT, and A and we vary which sections are involved in loss calculation. The six conditions are A, P+A, CoT, CoT+A, P+CoT, and P+CoT+A (full-length with all sections included). Following the training setup of previous work (Mukherjee et al., 2023; Mitra et al., 2023; Jiang et al., 2023; Gu et al., 2024; Shi et al., 2024; Shen et al., 2025; Li et al., 2025b), we apply a loss mask to all tokens outside the supervised region before computing the objective, while still feeding the entire P+CoT+A sequence through the model in the forward pass. If a section is not included, tokens within that section are conditioned but not involved in loss computation.
@@
Sequence-Length Scaling Behavior. First, we explore how performance scales with the number of tokens trained. Unlike loss masking, which still incurs forward-pass costs, we examine the performance when tokens are removed entirely. We introduce a dynamic truncation protocol, Lead-Span Proportion (LSP). For each example with tokenized inputs $\mathbf{x}$ and labels $\mathbf{y}$, LSP retains the first $p\in\{0.1,0.2,\dots,1.0\}$ fraction of tokens from the concatenation of $\mathbf{x}$ and $\mathbf{y}$, and removes the rightmost (1p)(1-p) fraction before training. The truncated tokens, therefore, never enter the model: they are neither seen in the forward pass nor included in the loss. This protocol preserves example-specific variability in kept length, avoids exclusion of answer tokens that can arise under a global max-length cutoff, and retains diversity of sequence lengths.
Budget Location Ablation. Second, to determine where the token budget should be spent, we fix the training budget at $50\%$ of the tokens (a choice guided by performance saturation, compute efficiency, and experimental symmetry, as detailed in Section 4.2). We compare two conditions: Left 50%, which keeps the first half of tokens and drops the rest, and Right 50%, which keeps the final half and drops the beginning. This ablation keeps the compute budget fixed while shifting the kept segment of tokens, enabling a clean test of whether early tokens carry richer learning signal than later ones. Better performance in the Left 50% setting would be consistent with the hypothesis that useful signal concentrates early in the trajectory.
@@ 4.1 Section-wise Supervision (RQ1)
Sub-optimality of full sequence distillation.
As can be seen in Figure 1, a consistent pattern emerges across benchmarks and teacher-student pairs: settings that supervise with CoT consistently outperform those that do not. On both benchmarks, answer-only supervision (A) yields the lowest accuracy, and extending the loss to the prompt (P+A) brings only modest gains. Once CoT tokens receive supervision, whether alone (CoT), paired with the answer (CoT+A), paired with the prompt (P+CoT), or combined with both (P+CoT+A), performance improves substantially, with only minor differences among these CoT-inclusive variants.
@@
Overall, the results indicate that supervising the rationale tokens is crucial, while spreading loss to the prompt and answer offers only modest, sometimes negligible, additional benefits. Thus, full-token supervision does not reliably outperform more selective objectives that focus on the reasoning process itself.
CoT implicitly encompasses the prompt and answer sections.
To understand the limited gains from supervising P and A, we manually examined the data and observed that the CoT often encapsulates information from both the prompt and the answer as illustrated in Figure 2. Typically, the CoT begins by paraphrasing the question to contextualize the problem, then proceeds with reflection-based reasoning steps before deriving the final answer. The answer section typically summarizes the CoT and reiterates the final answer near the end.
@@
Linguistic analysis.
To verify whether P and A are commonly covered by the CoT in our training datasets, we conduct a linguistic analysis on 100 randomly sampled training examples using GPT-5.1 as a judge. We prompt the model to verify that: (1) all substantive information in P and A is already present in the CoT; and (2) P and A do not introduce any new substantive statements absent from the CoT. This ensures that, semantically, the CoT entails both P and A. We also checked whether the final answer in A is equivalent to the final answer derived in CoT.
@@
Figure 5: First half of tokens tends to contain denser useful information than the second half, in line with stronger performance when training on the left $50\%$ of tokens. Complete examples, each for one of the five datasets, are included in Appendix A.1.1.
@@
Marginal benefits of longer reasoning sequences.
Figure 3 shows that across all distillation settings, LSP exhibits similar scaling behavior of the student model performance. They show nearly monotonic accuracy improvement as LSP increases and then saturates. The flattening near the right edge suggests those trimmed tokens are not required for effective distillation, implying substantial compute savings with minimal performance drop.
Role of model size.
The 32B-8B setting is consistently the top performer, and the ranking is largely determined by student size: 8B > 4B > 1.7B > 0.6B. For a fixed student size, curves from different teachers track closely. For the 4B student model, it performs comparably and sometimes better when distilling from 8B rather than 32B model, which echoes the findings of Li et al. (2025c) that smaller models do not necessarily learn better from much larger teachers. For small students (0.6B and 1.7B), training with LSP does not reliably beat the baselines. Their curves hew close to the dotted baselines across $p$, suggesting limited capacity to absorb additional supervision. The benefits of LSP mainly materialize once the student has enough scale (4B and 8B). Therefore, in later experiments, these two small students will not be involved.
@@
Deciding the most cost-effective budget. To anchor the token budget allocation, we select $p=0.5$ as our target. This decision is driven by three factors: performance saturation, computational efficiency, and experimental rigor. First, we identified the most cost-effective point (“knee”) in the accuracy-vs-LSP curves with  Satopaa et al. (2011) [^14] https://github.com/arvkevi/kneed, as shown in Table 3. We observe that across all settings, the knee consistently occurs at $p\leq 0.6$, with the majority at or below $p=0.5$. Second, as detailed in Section 4.3, $p=0.5$ serves as a critical boundary before computational costs begin to rise drastically. Third, selecting $p=0.5$ enables a clean, symmetrical ablation between the Left 50% and Right 50% of the sequence. A budget of $p>0.5$ would necessitate overlapping tokens. As seen in Table 4, across all 30 settings (3 teacher-student pairs, 5 datasets, and 2 benchmarks), training with the first half of tokens retains on average $\approx 91\%$ of the accuracy when compared to training with all tokens, with retention being as high as $\approx 99\%$ for AIME25 with 8B-4B setup.
OpenThoughts Bespoke Synthetic-1 Nemotron SkyT1
Pairs 24 25 24 25 24 25 24 25 24 25
32B-4B 96.21 97.40 97.98 93.28 79.28 91.77 96.56 90.71 87.41 89.65
8B-4B 98.07 98.92 85.79 87.62 82.54 92.12 91.50 91.62 75.23 89.01
32B-8B 95.96 97.63 92.43 87.16 95.23 97.06 92.21 96.63 80.12 82.66
@@
Table 4: The performance retention rate (%) training with the first half of the tokens of each example in five datasets, compared to training with full-length examples. Performance is evaluated on AIME24 (“24”) and AIME25 (“25”). See Appendix A.5.1 for the raw values used to compute the retention rate.
@@
Figure 6: GPU memory usage, time spent for each training step, and FLOPs as functions of Lead-Span Proportion (LSP). Due to the linear behavior of flops, not all but only selected LSPs ($0.5,1.0)$ for some teacher-student pairs are profiled.
@@
Why Early Tokens Suffice? The fact that the first $50\%$ of tokens yields similar performance compared to full-length raises a mechanistic question: does the early segment already contain the key reasoning signal? To probe this, we analyze where the first derivation of the correct answer appears inside the CoT, treating it as a proxy for the earliest point at which the reasoning becomes sufficient. For each of $100$ randomly sampled training examples, we prompt GPT-5.1 to identify the first derivation of the correct answer, and then compute its relative position within the sequence.
Across datasets, the first correct derivation tends to occur near the midpoint of the whole sequence; see Appendix A.2 for details. We illustrate this phenomenon in Figure 5. This near-midpoint tendency aligns with the accuracy–vs–LSP curves exhibiting knees at $p\leq 0.6$ (Table 3) and with the strong performance retention at $p=0.5$ (Table 4). Together, these results suggest that an early budget captures most of the learnable signal for distillation.
@@
Qualitative analysis of the content preceding the first derivation shows that this segment contains question understanding, initial ideation, several rounds of self-reflections, and substantial progress toward the final answer. In many examples, the remaining tokens are then largely dominated by verification, minor algebraic clean-up, and re-statements of an already-correct answer, which appear less informative for distillation than the earlier reasoning steps.
@@
We operationalize self-reflection, a key contributor to the capabilities of reasoning models Muennighoff and others (2025); Li et al. (2025a), by counting occurrences of lexical cues (e.g., “wait”, and related variants; see full list in Appendix A.2.1) that appear before the first correct derivation point. On average, we observe $2.2$ (OpenThoughts), $2.4$ (Bespoke), $2.8$ (Synthetic-1), $5.1$ (Nemotron), and $1.9$ (SkyT1) self-reflection cues before the first correct derivation.
@@
Ablation Studies Having established that the left $50\%$ often encapsulates the core reasoning process, we use the Left vs. Right ablation to verify that the efficiency gains are specific to this self-contained initial segment, rather than a general property of reducing the token count. We compare training with the Left $50\%$ against the Right $50\%$. As shown in Figure 4, training with the left $50\%$ of the sequence consistently and significantly outperforms training with the right $50\%$. This is further supported by our loss analysis (Appendix A.3), which reveals that the loss on the left 50% is comparable to the full sequence, whereas the loss on the right 50% is relatively higher, indicating that the initial segment is more amenable to learning.
@@ 4.3 Compute Efficiency
Computationally, our $50\%$ training protocol yields substantial resource savings, as illustrated in Figure 6. We observe a critical inflection point at LSP =0.5=0.5 for both GPU memory usage and training step time, beyond which costs escalate rapidly due to the quadratic complexity of attention mechanisms.
@@
Training Speed. Training step time follows a similar trajectory. While larger students generally require more time, we also observe that teacher size imposes a distinct overhead; for the same 4B student, distilling from a 32B teacher is consistently slower than from an 8B teacher. By capping the sequence at $50\%$, we achieve a roughly $2\times$ reduction in both training step time and FLOPs, confirming that $p=0.5$ is a practical sweet spot for resource-constrained distillation.
@@ 5 Conclusion
In this work, we demonstrate that the reasoning trace itself is the dominant carrier of supervisory signal in knowledge distillation. By systematically varying the supervised sections of training examples, we found that models trained on the reasoning trace alone achieve performance comparable to full-sequence supervision. Moreover, our investigation into token budget allocation reveals that critical reasoning behaviors are concentrated in early tokens; truncating examples to just their first $50\%$ of tokens retains $\approx 91\%$ of downstream accuracy on math tasks while reducing training time and memory usage by $\approx 50\%$. This establishes sequence truncation as a powerful efficiency lever, positioning partial-sequence training as a cost-effective strategy for scaling reasoning distillation.
Limitations
Our study focuses on mathematical reasoning tasks. While mathematics provides a rigorous testbed for reasoning with verifiable ground truth, extending these findings to other reasoning domains (e.g., commonsense reasoning, and code generation) remains an avenue for future work.
We conduct experiments within the Qwen3 model family, using Qwen3-32B and Qwen3-8B as teachers and Qwen3-8B/4B/1.7B/0.6B as students. This controlled setup isolates the effects of sequence truncation from confounding factors introduced by architectural heterogeneity across model families. Investigating cross-family distillation (e.g., Qwen to Gemma) represents a natural extension.
Ethical Considerations
This work investigates efficient training protocols for knowledge distillation. All datasets and models used are publicly available and were created by prior work for research purposes.
Our findings on sequence truncation have potential positive implications for democratizing access to reasoning model training, as they substantially reduce computational requirements. By demonstrating that comparable performance can be achieved with approximately half the training compute and memory, this work may lower barriers for researchers and practitioners with limited computational resources.
We acknowledge that improvements in reasoning model distillation could, in principle, be applied to both beneficial and harmful use cases. However, the efficiency techniques we propose do not fundamentally alter model capabilities. They reduce training costs for existing distillation paradigms. We believe the benefits of making reasoning distillation more accessible outweigh potential risks.
@@ A.1.1 Full Training Examples
Figure 9: A complete training example from Synthetic-1 sampled for linguistic analysis with three sections divided for illustration purposes. Red boxes indicate self-reflection keywords. Blue underline is the first derivation of correct answer labeled by GPT-5.1.
Figure 10: A complete training example from Nemotron sampled for linguistic analysis with three sections divided for illustration purposes. Red boxes indicate self-reflection keywords. Blue underline is the first derivation of correct answer labeled by GPT-5.1.
@@ A.1.2 Training Dataset Token Count Distribution
Figure 14: Token count distributions for prompts, chain-of-thought, answers, and full sequences in the sampled 16k training examples from Synthetic-1 dataset.
Figure 15: Token count distributions for prompts, chain-of-thought, answers, and full sequences in the sampled 16k training examples from Llama-Nemotron-Post-Training-Dataset dataset.
@@ A.2.2 Results
We present the results of our linguistic analysis on all five datasets in Table 6.
@@
P $\subseteq$ CoT A $\subseteq$ CoT Ans Match CoT Pos. Full Seq Pos. Self-Refl.
OpenThoughts 99 89 100 50.5% 43.3% 2.23
Bespoke 97 93 99 47.5% 42.0% 2.42
Synthetic-1 96 94 96 46.6% 40.7% 2.76
Nemotron 99 97 99 61.5% 55.9% 5.13
SkyT1 90 73 91 66.8% 54.0% 1.91
Table 6: Linguistic analysis results on 100 sampled examples per dataset. Columns indicate the percentage of examples where Prompt (P) and Answer (A) are fully covered by the CoT, whether the final answer matches, the average relative token position of the first correct answer derivation within CoT segment (CoT Pos.), the average relative token position of the first correct answer derivation within full sequence (Full Seq Pos.), and the average count of self-reflection keywords preceding that derivation (Self-Refl.)
Semantic entailment.
Across the four DeepSeek-R1-synthesized datasets (OpenThoughts, Bespoke, Synthetic-1, and Nemotron), the Prompt is almost always covered by the CoT (P $\subseteq$ CoT: 96–99%), and the Answer is also largely redundant given the CoT (A $\subseteq$ CoT: 89–97%). Moreover, the final answer derived in the CoT nearly always matches the final boxed answer (Ans Match: 96–100%). This supports the qualitative pattern described in §4.1: CoT is the dominant carrier of learnable signal, while adding loss on P and A often provides only marginal extra information.
First correct derivation position.
The “CoT Pos.” column reports the average relative token position within the CoT segment of the first derivation of the correct answer. The “Full Seq Pos.” column maps the same point to the full sequence by accounting for the prompt prefix length. Using the average section token shares in Table 2, we compute
$\text{Full Seq Pos.}=Share_{\text{prompt}}+(\text{CoT Pos.})\times Share_{\text{cot}},$
where Full Seq Pos. denotes the relative position in the full sequence.
Self-reflection
The “Self-Refl.” column indicates that self-correction language appears before the first correct derivation point, suggesting that early tokens often include iterative checking and revision. Notably, Nemotron exhibits the highest density of such cues (5.13 per example), consistent with more explicit verification-style traces, whereas SkyT1 shows fewer self-reflection cues on average (1.91).
@@
SkyT1 shows lower redundancy than the DeepSeek-R1 datasets: P $\subseteq$ CoT drops to 90% and A $\subseteq$ CoT drops to 73%, and the final answer agreement is also lower (Ans Match: 91%). This indicates that, relative to DeepSeek-R1-style traces, SkyT1 is more likely to place substantive content outside the canonical CoT span and/or to introduce additional reasoning after the CoT segment. This is consistent with the slightly lower retention observed when training with truncated sequences on SkyT1: if later tokens are more likely to contain corrections or new derivations, truncation becomes more lossy. We observe an example where the CoT derives a final answer (e.g., 47) near the end of the CoT span, but the Answer section continues the derivation and reaches a different final answer (e.g., 54):